# Beyond Slater’s Condition in Online CMDPs with Stochastic and Adversarial Constraints

### v1

@@ B.3 Results on the Weights
$\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau,i}(x,a))\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}$
$\mkern 90.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}+\|q_{\tau}-\widehat{q}_{\tau}\|_{1}$ (13)
$\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau,i}(x,a))\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}$
$\mkern 90.0mu+2L|X|\sqrt{2(t_{2}-t_{1})\ln\frac{2L}{\delta}}+3L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}}$ (14)
$=\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{(\widehat{g}_{\tau+1,i}(x,a)-\widehat{g}_{\tau,i}(x,a))}{\beta_{\tau,i}(x,a)}$
$\mkern 90.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}+2L|X|\sqrt{2(t_{2}-t_{1})\ln\frac{2L}{\delta}}$
$\mkern 90.0mu+3L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}$ (15)
$\leq\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{(\widehat{g}_{\tau+1,i}(x,a)-\widehat{g}_{\tau,i}(x,a))}{\beta_{\tau,i}(x,a)}$
$\mkern 90.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
where Equation (12) is due to the fact that $\widehat{q}_{\tau}\in\widehat{\Delta}_{\tau}(\mathcal{P}_{\tau})$, Inequality (13) holds by Lemma C.3, from which we have, with probability $1-\delta$:
$\sum_{\tau=t_{1}}^{t_{2}}(g_{\tau,i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau,i}(x,a))\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}.$
Inequality (14) follows from Lemma B.3 of [Rosenberg and Mansour, 2019b], with probability at least $1-2\delta$—notice that, all the results mentioned above can be trivially extended to hold in the interval $t_{1},t_{2}]$—. Equation (15) holds by the definition of the update:
$\widehat{g}_{\tau,i}(x,a)=\left(1-\beta_{\tau,i}(x,a)\right)\widehat{g}_{\tau,i}(x,a)+\beta_{\tau,i}(x,a)g_{\tau,i}(x,a),\quad\forall(x,a)\ \text{such that}\ \mathbb{I}_{\tau}\{x,a\}=1.$
A final Union Bound concludes the proof. ∎
@@
$V_{[t_{1},t_{2}],i}\leq\sum_{x\in X,a\in A}\frac{2}{\beta_{\ell(x,a,[t_{1},t_{2}]),i}(x,a)}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
@@
$\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau+1,i}(x,a)-\widehat{g}_{\tau,i}(x,a)\right)\leq\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{t_{2},i}(x,a)}$
Fixing a $(x,a)\in X\times A$ and defining $h=|\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]|$ as the number of times the pair $(x,a)$ is visited in the interval $t_{1},t_{2}]$,let $\tau(j)$ be the rounds in which the pair $(x,a)$ is visited the $j^{\text{th}}$ time in $t_{1},t_{2}]$. Then we have:
$\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau+1,i}(x,a)-\widehat{g}_{\tau,i}(x,a)\right)$
$\mkern 50.0mu=\sum_{j\in[h]}\frac{1}{\beta_{\tau(j),i}(x,a)}\left(\widehat{g}_{\tau(j+1),i}(x,a)-\widehat{g}_{\tau(j),i}(x,a)\right)$
$\mkern 50.0mu=\sum_{j\in[h-1]}\left(\frac{1}{\beta_{\tau(j),i}(x,a)}\left(\widehat{g}_{\tau(j+1),i}(x,a)-\widehat{g}_{\tau(j),i}(x,a)\right)\right)$
$\mkern 300.0mu+\frac{1}{\beta_{\tau(h),i}(x,a)}\left(\widehat{g}_{\tau(h)+1,i}(x,a)-\widehat{g}_{\tau(h),i}(x,a)\right)$
$\mkern 50.0mu\leq\sum_{j\in[h-1]}\left(\frac{1}{\beta_{\tau(j+1),i}(x,a)}\widehat{g}_{\tau(j+1),i}(x,a)-\frac{1}{\beta_{\tau(j),i}(x,a)}\widehat{g}_{\tau(j),i}(x,a)\right)$
$\mkern 300.0mu+\frac{1}{\beta_{\tau(h),i}(x,a)}\left(\widehat{g}_{\tau(h)+1,i}(x,a)-\widehat{g}_{\tau(h),i}(x,a)\right)$ (16)
$\mkern 50.0mu=\frac{1}{\beta_{\tau(h),i}(x,a)}\widehat{g}_{\tau(h)+1,i}(x,a)-\frac{1}{\beta_{\tau(1),i}(x,a)}\widehat{g}_{\tau(1),i}(x,a)$ (17)
@@
where Inequality (16) and Inequality (18) follow from the hypothesis that the learning rates are decreasing in the interval and Equation (17) follows from evaluating the telescoping sum. ∎
@@ B.4 Concentration Results
Given $c>0$, $\alpha\in(0,1)$, $t\in[T]$ and $\delta\in(0,1)$, let $b_{t}(x,a)=\frac{c}{N_{t}(x,a)^{\alpha}}$ for all $(x,a)\in X\times A$. Then, with probability $1-3\delta$ it holds:
$\sum_{\tau=1}^{t}b_{\tau}^{\top}\widehat{q}_{\tau}\leq\frac{c}{1-\alpha}|X|^{\alpha}|A|^{\alpha}L^{1-\alpha}t^{1-\alpha}+7L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$
@@
$\sum_{\tau=1}^{t}b_{\tau}^{\top}\widehat{q}_{\tau}$ $=\sum_{\tau=1}^{t}b_{\tau}^{\top}\widehat{q}_{\tau}+\sum_{\tau=1}^{t}b_{\tau}^{\top}q_{\tau}-\sum_{\tau=1}^{t}b_{\tau}^{\top}q_{\tau}$
$\leq\|q_{\tau}-\widehat{q}_{\tau}\|_{1}+\sum_{\tau=1}^{t}b_{\tau}^{\top}q_{\tau}$
$\leq 2L|X|\sqrt{2T\ln\frac{2L}{\delta}}+3L|X|\sqrt{2T|A|\ln\frac{2T|X||A|}{\delta}}+\sum_{\tau=1}^{t}b_{\tau}^{\top}q_{\tau}$ (20)
@@
$\mkern 200.0mu+2L\sqrt{2T\ln\frac{1}{\delta}}+\sum_{\tau=1}^{t}\sum_{x\in X,a\in A}b_{\tau}(x,a)\mathbb{I}_{\tau}\{x,a\}$ (21)
@@
$\mkern 200.0mu+2L\sqrt{2T\ln\frac{1}{\delta}}+\frac{c}{1-\alpha}|X|^{\alpha}|A|^{\alpha}L^{1-\alpha}t^{1-\alpha}$ (22)
@@ B.5 Violation Bound
exploiting the fact that $N_{t}(x,a)\leq t$ for all $t\in[T]$. Therefore, by Corollary B.5 we can write:
$V_{[\underline{t},\bar{t}],i}$ $\leq\frac{2}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}\sum_{x\in X,a\in A}\sqrt{N_{\bar{t}}(x,a)}+\sum_{\tau=\underline{t}}^{\bar{t}}b_{\tau}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$
$\leq\frac{2\sqrt{|X||A|L\bar{t}}}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}+\sum_{\tau=\underline{t}}^{\bar{t}}b_{\tau}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$ (23)
$\leq\frac{2\sqrt{|X||A|L\bar{t}}}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}+2\sqrt{2|X||A|Lt\ln\left(\frac{2T^{2}|X||A|}{\delta}\right)}$
$\mkern 300.0mu+14L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$ (24)
@@ B.6 Towards the Regret Bound in the Stochastic Setting
$\leq\sum_{\tau=1}^{t}\sum_{x\in X,a\in A}\bar{g}_{i}(x,a)\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2t\ln\frac{1}{\delta}}$
$\mkern 100.0mu+2L|X|\sqrt{2t\ln\frac{2L}{\delta}}+3L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$ (27)
@@
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau,i})^{\top}\widehat{q}_{\tau}+\sum_{\tau=1}^{t-1}b_{\tau}^{\top}\widehat{q}_{\tau}+g_{t,i}^{\top}q_{t}+2L\sqrt{2t\ln\frac{1}{\delta}}+5L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}$ (30)
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+g_{t,i}^{\top}q_{t}$
$\mkern 100.0mu+2L\sqrt{2t\ln\frac{1}{\delta}}+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}\quad$ (31)
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 100.0mu+2L\sqrt{2t\ln\frac{1}{\delta}}+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}$
$\leq\sum_{\tau=1}^{t-1}(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 100.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}\quad$ (32)
$\leq\sum_{\tau=1}^{t-1}\sum_{x\in X,a\in A}\left(\bar{g}_{i}(x,a)-\widehat{g}_{\tau,i}(x,a)\right)\mathbb{I}_{\tau}\{x,a\}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 100.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}\quad$ (33)
$\leq\sqrt{2\ln\frac{2m|X||A|T}{\delta}}\sum_{x\in X,a\in A}\sum_{\tau=1}^{t-1}\frac{1}{\sqrt{N_{\tau}(x,a)}}\mathbb{I}_{\tau}\{x,a\}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 100.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}$
$\leq 2\sqrt{2|X||A|t\ln\frac{2m|X||A|T}{\delta}}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 100.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}$
@@
where Inequality (28) holds by Lemma C.3 with probability $1-\delta$, Inequality (29) holds by Lemma B.8 with probability at least $1-2\delta$, Inequality (30) holds because $\widehat{q}_{\tau}\in\widehat{\Delta}_{\tau}(\mathcal{P}_{\tau})$, Inequality (31) holds by Lemma B.6 with probability at least $1-3\delta$, taking $\alpha=\frac{1}{2}$ and $c=\sqrt{2\ln\left(\frac{2|X||A|T}{\delta}\right)}$, Inequality (32) holds by Lemma C.4 with probability at least $1-\delta$, Inequality (33) holds by Lemma B.9 with probability at least $1-3\delta$.

### v2

@@ B.3 Results on the Weights
$\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a))\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}$
$\mkern 30.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+\|q_{\tau}-\widehat{q}_{\tau}\|_{1}$ (10)
$\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a))\mathbb{I}_{\tau}\{x,a\}$
$\mkern 30.0mu+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}$
$\mkern 30.0mu+2L|X|\sqrt{2(t_{2}-t_{1})\ln\frac{2L}{\delta}}$
$\mkern 30.0mu+3L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}}$ (11)
$=\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{(\widehat{g}_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a))}{\beta_{\tau,i}(x,a)}$
$\mkern 30.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+2L|X|\sqrt{2(t_{2}-t_{1})\ln\frac{2L}{\delta}}$
$\mkern 30.0mu+3L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}}+2L\sqrt{2(t_{2}-t_{1})\ln\frac{1}{\delta}}$ (12)
$\leq\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{(\widehat{g}_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a))}{\beta_{\tau,i}(x,a)}$
$\mkern 30.0mu+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
where Equation (9) is due to the fact that $\widehat{q}_{\tau+1}\in\widehat{\Delta}_{\tau}(\mathcal{P}_{\tau})$, Inequality (10) holds by Lemma C.3, from which we have, with probability $1-\delta$:
$\sum_{\tau=t_{1}}^{t_{2}}(g_{\tau,i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}\leq\sum_{\tau=t_{1}}^{t_{2}}\sum_{x\in X,a\in A}(g_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a))\mathbb{I}_{\tau}\{x,a\}+L\sqrt{8(t_{2}-t_{1})\ln\frac{1}{\delta}}.$
Inequality (11) follows from Lemma B.3 of [Rosenberg and Mansour, 2019b], with probability at least $1-2\delta$—notice that all the results mentioned above can be trivially extended to hold in the interval $t_{1},t_{2}]$—. Equation (12) holds by the definition of the update:
$\widehat{g}_{\tau,i}(x,a)=\left(1-\beta_{\tau,i}(x,a)\right)\widehat{g}_{\tau-1,i}(x,a)+\beta_{\tau,i}(x,a)g_{\tau,i}(x,a),$
[…]
@@
$V_{[t_{1},t_{2}],i}\leq\sum_{x\in X,a\in A}\frac{2}{\beta_{\ell(x,a,[t_{1},t_{2}]),i}(x,a)}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
@@
$\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a)\right)\leq\sum_{x\in X,a\in A}\frac{2}{\beta_{t_{2},i}(x,a)}.$
Fixing a $(x,a)\in X\times A$ and defining $h=|\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]|$ as the number of times the pair $(x,a)$ is visited in the interval $t_{1},t_{2}]$,let $\tau(j)$ be the rounds in which the pair $(x,a)$ is visited the $j^{\text{th}}$ time in $t_{1},t_{2}]$.
Then we have:
$\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a)\right)$
$\mkern 50.0mu=\sum_{j\in[h]}\frac{1}{\beta_{\tau(j),i}(x,a)}\left(\widehat{g}_{\tau(j),i}(x,a)-\widehat{g}_{\tau(j-1),i}(x,a)\right)$
$\mkern 50.0mu=\sum_{j\in[h-1]}\left(\frac{1}{\beta_{\tau(j),i}(x,a)}\left(\widehat{g}_{\tau(j),i}(x,a)-\widehat{g}_{\tau(j-1),i}(x,a)\right)\right)$
$\mkern 100.0mu+\frac{1}{\beta_{\tau(h),i}(x,a)}\left(\widehat{g}_{\tau(h),i}(x,a)-\widehat{g}_{\tau(h-1),i}(x,a)\right)$
$\mkern 50.0mu\leq\sum_{j\in[h-1]}\left(\frac{1}{\beta_{\tau(j+1),i}(x,a)}\widehat{g}_{\tau(j),i}(x,a)-\frac{1}{\beta_{\tau(j),i}(x,a)}\widehat{g}_{\tau(j-1),i}(x,a)\right)$
$\mkern 100.0mu+\frac{1}{\beta_{\tau(h),i}(x,a)}\left(\widehat{g}_{\tau(h),i}(x,a)-\widehat{g}_{\tau(h-1),i}(x,a)\right)$ (13)
$\mkern 50.0mu=\frac{1}{\beta_{\tau(h),i}(x,a)}\widehat{g}_{\tau(h),i}(x,a)-\frac{1}{\beta_{\tau(1),i}(x,a)}\widehat{g}_{\tau(1)-1,i}(x,a)$ (14)
@@
where Inequality (13) and Inequality (15) follow from the hypothesis that the learning rates are decreasing in the interval, and Equation (14) follows from evaluating the telescoping sum. ∎
@@ B.4 Concentration Results
Given $c>0$, $\alpha\in(0,1)$, $t\in[T]$ and $\delta\in(0,1)$, let the bonus $b_{t}(x,a)=\max\left\{1,\frac{c}{N_{t}(x,a)^{\alpha}}\right\}$ for all $(x,a)\in X\times A$. Then, with probability $1-3\delta$ it holds:
$\sum_{\tau=1}^{t}b_{\tau-1}^{\top}\widehat{q}_{\tau}\leq\frac{c}{1-\alpha}|X|^{\alpha}|A|^{\alpha}L^{1-\alpha}t^{1-\alpha}+7L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$
@@
$\sum_{\tau=1}^{t}b_{\tau-1}^{\top}\widehat{q}_{\tau}$ $=\sum_{\tau=1}^{t}b_{\tau-1}^{\top}\widehat{q}_{\tau}+\sum_{\tau=1}^{t}b_{\tau-1}^{\top}q_{\tau}-\sum_{\tau=1}^{t}b_{\tau-1}^{\top}q_{\tau}$
$\leq\|q_{\tau}-\widehat{q}_{\tau}\|_{1}+\sum_{\tau=1}^{t}b_{\tau-1}^{\top}q_{\tau}$
$\leq 2L|X|\sqrt{2T\ln\frac{2L}{\delta}}+3L|X|\sqrt{2T|A|\ln\frac{2T|X||A|}{\delta}}+\sum_{\tau=1}^{t}b_{\tau-1}^{\top}q_{\tau}$ (17)
@@
$\mkern 20.0mu+2L\sqrt{2T\ln\frac{1}{\delta}}+\sum_{\tau=1}^{t}\sum_{x\in X,a\in A}b_{\tau-1}(x,a)\mathbb{I}_{\tau}\{x,a\}$ (18)
@@
$\mkern 20.0mu+2L\sqrt{2T\ln\frac{1}{\delta}}+L+\sum_{\tau=1}^{t}\sum_{x\in X,a\in A}b_{\tau}(x,a)\mathbb{I}_{\tau}\{x,a\}$
$\leq 2L|X|\sqrt{2T\ln\frac{2L}{\delta}}+3L|X|\sqrt{2T|A|\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 20.0mu+2L\sqrt{2T\ln\frac{1}{\delta}}+\frac{c}{1-\alpha}|X|^{\alpha}|A|^{\alpha}L^{1-\alpha}t^{1-\alpha}$ (19)
@@ B.5 Violation Bound
exploiting the fact that $N_{t}(x,a)\leq t$ for all $t\in[T]$.
Therefore, by Corollary B.5, since the constraints learning rates are decreasing in the interval, the following holds:
$V_{[\underline{t},\bar{t}],i}$ $\leq\frac{2}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}\sum_{x\in X,a\in A}\sqrt{N_{\bar{t}}(x,a)}+\sum_{\tau=\underline{t}}^{\bar{t}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$
$\leq\frac{2\sqrt{|X||A|L\bar{t}}}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}+\sum_{\tau=\underline{t}}^{\bar{t}}b_{\tau-1}^{\top}\widehat{q}_{\tau}+7L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$ (20)
$\leq\frac{2\sqrt{|X||A|L\bar{t}}}{21L\sqrt{|X||A|t\ln\left(\frac{T^{2}}{\delta}\right)}}+2\sqrt{2|X||A|Lt\ln\left(\frac{2T^{2}|X||A|}{\delta}\right)}+14L|X|\sqrt{2t|A|\ln\frac{2T^{2}|X||A|}{\delta}}$ (21)
@@ B.6 Towards the Regret Bound in the Stochastic Setting
$\leq\sum_{\tau=1}^{t}\sum_{x\in X,a\in A}\bar{g}_{i}(x,a)\mathbb{I}_{\tau}\{x,a\}+2L\sqrt{2t\ln\frac{1}{\delta}}+2L|X|\sqrt{2t\ln\frac{2L}{\delta}}+3L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$ (24)
@@
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau-1,i})^{\top}\widehat{q}_{\tau}+\sum_{\tau=1}^{t-1}b_{\tau-1}^{\top}\widehat{q}_{\tau}+g_{t,i}^{\top}q_{t}+2L\sqrt{2t\ln\frac{1}{\delta}}$
$\mkern 30.0mu+5L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}$ (27)
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau-1,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+g_{t,i}^{\top}q_{t}$
$\mkern 30.0mu+2L\sqrt{2t\ln\frac{1}{\delta}}+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}\quad$ (28)
$\leq\sum_{\tau=1}^{t-1}(g_{\tau,i}-\widehat{g}_{\tau-1,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 30.0mu+2L\sqrt{2t\ln\frac{1}{\delta}}+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}$
$\leq\sum_{\tau=1}^{t-1}(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}\widehat{q}_{\tau}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 30.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}\quad$ (29)
$\leq\sum_{\tau=1}^{t-1}\sum_{x\in X,a\in A}\left(\bar{g}_{i}(x,a)-\widehat{g}_{\tau-1,i}(x,a)\right)\mathbb{I}_{\tau}\{x,a\}$
$\mkern 30.0mu+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 30.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}\quad$ (30)
$\leq\sqrt{2\ln\frac{2m|X||A|T}{\delta}}\sum_{x\in X,a\in A}\sum_{\tau=1}^{t-1}\frac{1}{\sqrt{N_{\tau-1}(x,a)}}\mathbb{I}_{\tau}\{x,a\}$
$\mkern 30.0mu+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+L$
$\mkern 30.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}$
$\leq 2\sqrt{2|X||A|t\ln\frac{2m|X||A|T}{\delta}}+2\sqrt{2|X||A|Lt\ln\frac{2T|X||A|}{\delta}}+2L$
$\mkern 30.0mu+12L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}+4L\sqrt{2t\ln\frac{1}{\delta}}$
@@
where Inequality (25) holds by Lemma C.3 with probability $1-\delta$, Inequality (26) holds by Lemma B.8 with probability at least $1-2\delta$, Inequality (27) holds because $\widehat{q}_{\tau+1}\in\widehat{\Delta}_{\tau}(\mathcal{P}_{\tau})$, Inequality (28) holds by Lemma B.6 with probability at least $1-3\delta$, taking $\alpha=\frac{1}{2}$ and $c=\sqrt{2\ln\left(\frac{2|X||A|T}{\delta}\right)}$, Inequality (29) holds by Lemma C.4 with probability at least $1-\delta$, Inequality (30) holds by Lemma B.9 with probability at least $1-3\delta$.